OpenStudy (anonymous):

Solve for f(x)=g(x) when f(x)=-6/(x+3) and g(x)= 2/x

OpenStudy (anonymous):

-6/(x+3)=2/x solve the equation..

OpenStudy (anonymous):

\[\frac{-6}{x+3} = \frac{2}{x}\] \[-6x = 2(x+3)\]

OpenStudy (anonymous):

@jim_thompson5910 Can you walk me through this one please?

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

f(x)=g(x) -6/(x+3)=2/x -6x=2(x+3) -6x=2x+6 -6x-2x=6 -8x=6 x=6/(-8) x=-3/4 So the solution is x = -3/4

OpenStudy (anonymous):

Thank you so much! =)

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad I could help you out again :)